Jessica wants to accumulate $10,000 by the end of 6 years in a special bank account, which she had opened for this purpose. To achieve this goal, Jessica plans to deposit a fixed sum of money into the account at the end of the month over the 6-year period. If the bank pays interest at the rate of 6% per year compounded monthly, how much does she have to deposit each month into her account? (Round your answer to the nearest cent.) $
Answer #1

[(1+г) — 1 FV of Annuity %3D Р т Р- Рeriodic Pаутent г3 rate per period п 3 питber o ff periods

n = 6 * 12 = 72 months

r = 6%/12 = 0.50% (monthly)

10000 = P * [\frac{(1+0.005)^{72}-1}{0.005}]

10000 = P * 86.4089

P = $115.7